OpenStudy (anonymous):

equation of a conic section : Ellipse - foci (2,-4) & (5,-4).. Verticies (-1,-4) & (8,-4) ???

OpenStudy (bibby):

I think this site has what we want. Try matching the information to the equation of ellipse: Ax2 + Cy2 + Dx + Ey + F = 0

OpenStudy (bibby):

http://www.purplemath.com/modules/conics.htm

OpenStudy (anonymous):

okay thanks

OpenStudy (bibby):

First, we'll find the center using the midpoint formula. (-1,-4) & (8,-4) -> \[\large \frac{-1+8}{2},\frac{-4+-4}{2}=>(\frac{7}{2},-4)\]

OpenStudy (anonymous):

ooooookkk then what

OpenStudy (bibby):

I'm just taking this step by step following this problem. I have no idea what I'm doing

OpenStudy (anonymous):

weelll at least ur tyring thts more then anyone else is f'ing doing .... thnx

OpenStudy (anonymous):

cuz i was dead serious when i said i was in a group home and i need to graduate before i can go home

OpenStudy (bibby):

My wife left me. We all have our problems. Sadly, unless you're paying our lives don't revolve around you. The best we can do is help at our leisure

OpenStudy (shamil98):

tuf lyfe man i got u bruh aight les get it you need the general equation of an ellipse right?

OpenStudy (anonymous):

thanks buddy figured that out but unless ur persistant enough nothings gonna get done for u or anyone else for that matter

OpenStudy (anonymous):

yeah sham

OpenStudy (shamil98):

first, we need to see if the ellipse is at the origin of a plane

OpenStudy (shamil98):

so ill wolfram the coordinates sec.

OpenStudy (anonymous):

its the equation for a conic section idk if thts the same thing

OpenStudy (shamil98):

\[\large \frac{ x^2 }{ a^2 } + \frac{ y^2 }{ b^2 } = 1\] if its centered at the origin of a plane a is the radius along the x-axis b is the radius along the y-axis if its not then: \[\frac{ (x-h)^2 }{ a^2 } + \frac{ (y-k)^2 }{ b^2 } = 1\] a is the radius along the x-axis b is the radius along the y-axis h, k are the x,y coordinates of the ellipse's center try graphing the ellipse first with the foci/vertices

OpenStudy (anonymous):

ok gimme a sec.

OpenStudy (anonymous):

ok done .... now what

OpenStudy (bibby):

plug in the center I gave you into the second equation shamil gave you

OpenStudy (anonymous):

in english ...... plz

OpenStudy (shamil98):

plug in (7/2, -4) into the second equation i put up there ok? also use the image up there that bib put so u can find the values of a and b